As a foreign exchange trader, you see the following quotes for Canadian Dollars (CAD), U.S. Dollars (USD), and Mexican Pesos (MXN):
MXN 6.4390/CAD USD 0.7047/CAD MXN 8.7535/USD Is there an arbitrage opportunity? If so, how would you exploit it if you had 1 million USD? (11 marks)
Answer #1

Start with 1 million USD
Convert USD to CAD =1/0.7047 million CAD
Convert CAD to MXN=1/0.7047*6.4390 million MXN
Convert MXN to USD=1/0.7047*6.4390*1/8.7535=1.043836353

Profit=1.043836353-1=0.043836353 million USD or 43836.353 USD